OpenStudy (anonymous):

cos2 (x-y/2) - sin2 (x+y/2) = cosx.cosy

OpenStudy (anonymous):

you want to proove it ?!

OpenStudy (anonymous):

!??

OpenStudy (anonymous):

This? \[\cos^2(x-y/2)-\sin^2(x+y/2) = \cos(x) \cos(y)\] Or this? \[\cos^2(\frac{x-y}{2})-\sin^2(\frac{x+y}{2}) = \cos(x) \cos(y)\] And was there a question here? Are you solving for x, y, or what? The first equation does have two y solutions, a \(\pm\) constant

OpenStudy (anonymous):

Also given the level of difficulty with equations like this, please remember that on computers to show exponents you'll need to make use of the ^ symbol. i.e. x^2 = x\(^2\) x^(1/2) = x\(^{1/2}\) = \(\sqrt{\text{x}}\)

OpenStudy (anonymous):

cos2 = cos(2) \(\approx -0.4161468365\)